Alpha Ltd is considering a capital investment proposal where two alternatives, involving differing degrees of mechanization, are being considered. New machinery would cost Rs. 2,78,000 in option 1 and Rs. 8,05,000 in option 2. Both these investments would have a five year life, however it is anticipated that the scrap values after 5 years would be Rs. 28,000 and Rs. 1,50,000 respectively. Depreciation is provided on a straight line basis. Annual Cash Inflows, expected to be generated for option 1 is Rs. 1,00,000 and Rs. 2,50,000 for option 2. The cost of capital is 15%. You are required to calculate the following for each option: (i) The payback period (ii) The Accounting rate of return, based on average book value (iii) The Net Present Value (iv) The Internal Rate of Return. As a Manager which option you would prefer, give reasons for your preference.


1
Expert's answer
2021-05-27T16:11:01-0400

i) payback period option 1= 278000/100000= 2.7 years

Option 2= 805000/250000= 3.2 years

ii) ARR for option one = Average annual profit/ average investment where average annual profit is the total profit divided by 2 and average investment is total investment divided by 2

ARR option 1= 2.014

ARR option 2= 2.68

iii) Net present value for option 1=100000-0= 100,000

For option 2= 250,000- 0= 250,000

iii) IRR option 1= 23.4%

IRR option 2= 16.7%